OpenStudy (anonymous):

Simplify. –22 + (–5) ∙ 6

OpenStudy (anonymous):

try PEMDAS :o

OpenStudy (anonymous):

no i need help

OpenStudy (anonymous):

and it says simplify

OpenStudy (anonymous):

@TheSmartOne

OpenStudy (anonymous):

@triciaal

OpenStudy (anonymous):

@TuringTest

OpenStudy (anonymous):

so?

OpenStudy (anonymous):

did you try?

OpenStudy (anonymous):

PEMDAS?

OpenStudy (anonymous):

yep

OpenStudy (triciaal):

order of operations multiplication before addition and subtraction start with -5* 6 = then subtract 22

OpenStudy (anonymous):

ok so then it would be 30 then 30-22=?

OpenStudy (anonymous):

right?

OpenStudy (anonymous):

hello?

TheSmartOne (thesmartone):

No, it would become -22 - 30 = ?